# The $v^{1/3}_{3}/v^{1/2}_{2}$ ratio in PbAu collisions at $\sqrt{s_{\mathrm{NN}}}=$ 17.3 GeV

### v1

@@ Abstract
The Fourier harmonics, $v_{2}$ and $v_{3}$ of negative pions are measured at center-of-mass energy per nucleon pair of $\sqrt{s_{\mathrm{NN}}}$= 17.3 GeV around midrapidity by the CERES/NA45 experiment at the CERN SPS in 0–30% central PbAu collisions with a mean centrality of 5.5%. The analysis is performed in two centrality bins as a function of the transverse momentum $\mathrm{p_{\mathrm{T}}}$ from 0.05 GeV/$c$ to more than 2 GeV/$c$. This is the first measurement of the $v^{1/3}_{3}/v^{1/2}_{2}$ ratio as a function of transverse momentum at SPS energies. For $\mathrm{p_{\mathrm{T}}}$ above 0.5 GeV/$c$, the ratio is nearly flat in accordance with the hydrodynamic prediction and as previously observed by the ATLAS and ALICE experiments at the much higher LHC energies. The results are also compared with the SMASH-vHLLE hybrid model predictions.
@@ 1 Introduction
In this paper, we present the first result of negative pions $v^{1/3}_{3}/v^{1/2}_{2}$ ratio as a function of transverse momentum in PbAu collisions at the SPS at a nucleon-nucleon center-of-mass energy of $\sqrt{s_{\mathrm{NN}}}=$ 17.3 GeV. The analysis is based on previously published $v_{2}$ and $v_{3}$ data CERES:2012hbp ; CERESNA45:2016rra , where the $v_{2}$ data have been rebinned. The results are compared with the ATLAS and the ALICE measurements and show a very similar behavior. The results are also compared with predictions from the Simulating Many Accelerated Strongly-interacting Hadrons (SMASH) SMASH:2016zqf ; Karpenko:2015xea , a new hadron transport approach in heavy-ion collisions, combined with the vHLLE 3+1D viscous hydrosolver Karpenko:2013wva that simulates the QGP expansion. The paper is organized as follows. Some details concerning the experiment and the data sample are given in section 2. The data analysis and the obtained results are presented in section 3. A short summary is given in section 4.
@@ 3 Analysis and results
Figure 2 displays the HBT-corrected $v_{2}$ values of negative pions for two centrality classes in PbAu collisions at $\sqrt{s_{\mathrm{NN}}}=$ 17.3 GeV. The $v_{2}$ values of negative pions are taken from CERES:2012hbp and rebinned. In the same figure results from a hybrid model Karpenko:2015xea combining the vHLLE viscous hydrosolver Karpenko:2013wva with the hadronic transport model SMASH SMASH:2016zqf are plotted. The centrality ranges for the hydrodynamic calculation are comparable with the experimental ones. The model predictions are calculated for charged hadrons $-1<\eta<1$ that is close to the experimental acceptance. The value of the specific viscosity used in the vHLLE model is 0.15. For $p_{\mathrm{T}}$ above about 0.6 GeV/c, in both centrality classes the hydrodynamic prediction overestimates the data.
@@
In ideal hydrodynamics, at $p_{\mathrm{T}}\lesssim$ 2 GeV/c, the $v_{n}(p_{\mathrm{T}})$ behave as power-law functions of $p_{\mathrm{T}}$ Alver:2010dn ; Borghini:2005kd . This can be expressed as
@@
Figure 4 shows the ratio $v^{1/3}_{3}/v^{1/2}_{2}$ as a function of $p_{\mathrm{T}}$ for two centrality classes. The ratio is calculated for negative pions after suppression of the contribution of the HBT effect to the measured $v_{n}$ coefficients. The figure also shows the same ratio for PbPb collisions at $\sqrt{s_{\mathrm{NN}}}$= 2.76 TeV from the ATLAS ATLAS:2012at and ALICE ALICE:2018rtz experiment. The centrality classes are similar to those used by the CERES/NA45 experiment. The ATLAS and ALICE analyses have been performed on charged particles emitted within $|\eta|<2.5$ and $|\eta|<0.8$, respectively. There is a very good agreement between the ATLAS (blue squares) and ALICE (green squares) data. Despite the huge difference in incident energies between the top SPS energy and the LHC energy, for $p_{\mathrm{T}}$ above 0.5 GeV/c, the CERES/NA45 data are in a rather good agreement with the ATLAS and ALICE data, especially for the ‘top-central’ events. In the case of ‘mid-central’ events for $p_{\mathrm{T}}>$ 1.2 GeV/c, the CERES/NA45 results are somewhat above the ATLAS and ALICE data. Universally, there is a trend that the ratio for semicentral collisions is smaller. For the present data, there is a rising trend below 500 MeV/c, although not very significant within errors. The uncertainties are significantly larger for the current SPS data due to very small values of $v_{2,3}$. In the low-$p_{\mathrm{T}}$ range 0.05 – 0.5 GeV/c, where the HBT influence is largest, it may be that the HBT effect on the measured $v_{3}$ harmonic is not completely suppressed, thus producing larger values of the $v^{1/3}_{3}/v^{1/2}_{2}$ ratio. A $p_{\mathrm{T}}$-independent $v^{1/3}_{3}/v^{1/2}_{2}$ ratio is a hydrodynamic prediction Alver:2010dn ; Borghini:2005kd . The CERES/NA45 results imply that the system formed in PbAu collisions at the top SPS energy exhibits hydrodynamic behavior as already suggested in CERES:2012hbp .
@@
Figure 5 shows a comparison of the CERES $v^{1/3}_{3}/v^{1/2}_{2}$ data with the corresponding ratio calculated by the SMASH-vHLLE model. As expected, for $p_{\mathrm{T}}>$ 0.7 GeV/c, where the statistical uncertainties of the model are small enough, the model gives a flat ratio for both centrality classes. For $p_{\mathrm{T}}<$ 0.5 GeV/c, the statistical uncertainties of the model are very large. The SMASH-vHLLE model prediction tends to stay slightly below the experimental data. This is due to the fact that the model overestimates the experimentally measured $v_{2}$ somewhat for $p_{\mathrm{T}}>$ 0.7 GeV/c, while it reproduces $v_{3}$ harmonics, as shown in Figs. 2 and 3. These predictions of the hybrid model at SPS tend to be a little weaker than in data. In particular, the choice of initial state and $\eta/s$ value for hydrodynamic evolution is very important. This measurement could help to better constrain the initial conditions and specific viscosity.
@@ 4 Summary
The ratio $v^{1/3}_{3}/v^{1/2}_{2}$ of negative pions has been measured as a function of $p_{\mathrm{T}}$ using the data collected by the CERES/NA45 experiment for PbAu collisions at $\sqrt{s_{\mathrm{NN}}}=$ 17.3 GeV. This ratio, measured in two centrality classes, shows good agreement with the LHC results from the ATLAS and ALICE experiments. The SMASH-vHLLE hybrid model reproduce the experimentally measured $v_{3}(p_{\mathrm{T}})$, but somewhat overestimates the $v_{2}$ values at larger $p_{\mathrm{T}}$. Consequently, the model underestimates the measured ratio. These results shed some light on the dynamics of the system created in heavy-ion collisions at top SPS energy, and give additional constrain for the initial conditions and shear viscosity over entropy density.
@@ Acknowledgements.
The CERES collaboration acknowledges the good performance of the CERN PS and SPS accelerators as well as the support from the EST division. We are grateful for excellent support by the CERN IT division for the central data recording and data proccesing. This work was supported by GSI, Darmstadt, the German BMBF, the German VH-VI 146, the US DoE, the Israeli Science Foundation, and the MINERVA Foundation. We acknowledge the support by the Ministry of Education, Science and Technological Development of the Republic of Serbia throughout the theme 0102202. Z.P. acknowledges support by the State of Hesse within the Research Cluster ELEMENTS (Project ID 500/10.006). Computational resources for the SMASH-vHLLE hybrid modelling have been provided by the GreenCube at GSI.

### v2

@@ Abstract
The Fourier harmonics, $v_{2}$ and $v_{3}$ of negative pions are measured at center-of-mass energy per nucleon pair of $\sqrt{s_{\mathrm{NN}}}$= 17.3 GeV around midrapidity by the CERES/NA45 experiment at the CERN SPS in 0–30% central PbAu collisions with a mean centrality of 5.5%. The analysis is performed in two centrality bins as a function of the transverse momentum $\mathrm{p_{\mathrm{T}}}$ from 0.05 GeV/$c$ to more than 2 GeV/$c$. This is the first measurement of the $v^{1/3}_{3}/v^{1/2}_{2}$ ratio as a function of transverse momentum at SPS energies, that reveals, independently of the hydrodynamic models, hydrodynamic behavior of the formed system. For $\mathrm{p_{\mathrm{T}}}$ above 0.5 GeV/$c$, the ratio is nearly flat in accordance with the hydrodynamic prediction and as previously observed by the ATLAS and ALICE experiments at the much higher LHC energies. The results are also compared with the SMASH-vHLLE hybrid model predictions.
Keywords:
Fourier harmonics hydrodynamics SPS LHC heavy-ion
pacs:
25.75.Ld
journal: Eur. Phys. J. C
@@ 1 Introduction
In this paper, we present the first result of negative pions $v^{1/3}_{3}/v^{1/2}_{2}$ ratio as a function of transverse momentum in PbAu collisions at the SPS at a nucleon-nucleon center-of-mass energy of $\sqrt{s_{\mathrm{NN}}}=$ 17.3 GeV. The analysis is based on previously published $v_{2}$ and $v_{3}$ data CERES:2012hbp ; CERESNA45:2016rra , where the $v_{2}$ data have been rebinned. This reanalysis of the data provides a novel way of highlighting similarities with the flow phenomena seen at RHIC and LHC. The obtained results reveal, independently of the hydrodynamic models, a new interpretation of the SPS data which indicates hydrodynamic behavior of the system formed in PbAu collisions already at top SPS energy. The results are compared with the ATLAS and the ALICE measurements and show a very similar behavior. The results are also compared with predictions from the Simulating Many Accelerated Strongly-interacting Hadrons (SMASH) SMASH:2016zqf ; Karpenko:2015xea , a new hadron transport approach in heavy-ion collisions, combined with the vHLLE 3+1D viscous hydrosolver Karpenko:2013wva that simulates the QGP expansion. The paper is organized as follows. Some details concerning the experiment and the data sample are given in section 2. The data analysis and the obtained results are presented in section 3. A short summary is given in section 4.
@@ 3 Analysis and results
Figure 2 displays the HBT-corrected $v_{2}$ values of negative pions for two centrality classes in PbAu collisions at $\sqrt{s_{\mathrm{NN}}}=$ 17.3 GeV. The $v_{2}$ values of negative pions are taken from CERES:2012hbp and rebinned. In the same figure results from a hybrid model Karpenko:2015xea combining the vHLLE viscous hydrosolver Karpenko:2013wva with the hadronic transport model SMASH SMASH:2016zqf are plotted. The centrality ranges for the hydrodynamic calculation are comparable with the experimental ones. The model predictions are calculated for charged hadrons $-1<\eta<1$ that is close to the experimental acceptance. The value of the specific viscosity used in the vHLLE model is 0.15. For $p_{\mathrm{T}}$ above about 0.6 GeV/c, in both centrality classes the hydrodynamic prediction overestimates the data somewhat.
@@
In the hydrodynamics, at $p_{\mathrm{T}}\lesssim$ 2 GeV/c, the $v_{n}(p_{\mathrm{T}})$ behave as power-law functions of $p_{\mathrm{T}}$ Alver:2010dn ; Borghini:2005kd . This can be expressed as
@@
Figure 4 shows the ratio $v^{1/3}_{3}/v^{1/2}_{2}$ as a function of $p_{\mathrm{T}}$ for two centrality classes. The ratio is calculated for negative pions after suppression of the contribution of the HBT effect to the measured $v_{n}$ coefficients. The figure also shows the same ratio for PbPb collisions at $\sqrt{s_{\mathrm{NN}}}$= 2.76 TeV from the ATLAS ATLAS:2012at and ALICE ALICE:2018rtz experiment. The centrality classes are similar to those used by the CERES/NA45 experiment. The ATLAS and ALICE analyses have been performed on charged particles emitted within $|\eta|<2.5$ and $|\eta|<0.8$, respectively. There is a very good agreement between the ATLAS (blue squares) and ALICE (green squares) data. Despite the huge difference in incident energies between the top SPS energy and the LHC energy, for $p_{\mathrm{T}}$ above 0.5 GeV/c, the CERES/NA45 data are in a rather good agreement with the ATLAS and ALICE data, especially for the ‘top-central’ events. In the case of ‘mid-central’ events for $p_{\mathrm{T}}>$ 1.2 GeV/c, the CERES/NA45 results are somewhat above the ATLAS and ALICE data. Universally, there is a trend that the ratio for semicentral collisions is smaller. For the present data, there is a rising trend below 500 MeV/c, although not very significant within errors. The uncertainties are significantly larger for the current SPS data due to very small values of $v_{2,3}$ at the lower energy. In the low-$p_{\mathrm{T}}$ range 0.05 – 0.5 GeV/c, where the HBT influence is largest, it may be that the HBT effect on the measured $v_{3}$ harmonic is not completely suppressed, thus producing larger values of the $v^{1/3}_{3}/v^{1/2}_{2}$ ratio. A $p_{\mathrm{T}}$-independent $v^{1/3}_{3}/v^{1/2}_{2}$ ratio is a hydrodynamic prediction Alver:2010dn ; Borghini:2005kd . The CERES/NA45 results imply that the system formed in PbAu collisions at the top SPS energy exhibits hydrodynamic behavior as already suggested in CERES:2012hbp .
@@
Figure 5 shows a comparison of the CERES $v^{1/3}_{3}/v^{1/2}_{2}$ data with the corresponding ratio calculated by the SMASH-vHLLE model. As expected, for $p_{\mathrm{T}}>$ 0.7 GeV/c, where the statistical uncertainties of the model are small enough, the model gives a flat ratio for both centrality classes. For $p_{\mathrm{T}}<$ 0.5 GeV/c, the statistical uncertainties of the model are very large. The SMASH-vHLLE model prediction tends to stay slightly below the experimental data. This is due to the fact that the model overestimates the experimentally measured $v_{2}$ somewhat for $p_{\mathrm{T}}>$ 0.7 GeV/c, while it reproduces $v_{3}$ harmonics, as shown in Figs. 2 and 3.
This possibility was already discussed in our first publication of elliptic flow coefficients for pions where data were compared to ideal hydrodynamics calculations CERESNA45:2003pat as well as in CERES:2012hbp . Then we found that data could be well decsribed using ideal hydrodynamics, and a freeze-out temperature of 160 MeV, while for 120 MeV the flow was overpredicted for semicentral data. At the time we concluded that a supression mechanism should be at work not present in ideal hydrodynamic calculations. This is confirmed by the rather good description of the data presented in the current publication with a viscous hydrodynamics calculation. Albeit the comparison presented here is to results of a more recent hydrodynamics code, the supression is in magnitude entirely consistent with the value of $\eta/s=$ 0.15 used here.
These predictions of the hybrid model at SPS tend to be a little below data. In particular, the choice of initial state and $\eta/s$ value for hydrodynamic evolution is very important. This measurement could help to better constrain the initial conditions and specific viscosity.
@@ 4 Summary
The ratio $v^{1/3}_{3}/v^{1/2}_{2}$ of negative pions has been measured as a function of $p_{\mathrm{T}}$ using the data collected by the CERES/NA45 experiment for PbAu collisions at $\sqrt{s_{\mathrm{NN}}}=$ 17.3 GeV. This ratio, measured in two centrality classes, shows rather similar as the LHC results from the ATLAS and ALICE experiments. The SMASH-vHLLE hybrid model calculations reasonably well represent the data. The results give a new interpretation of the SPS data at its top energy which indicates hydrodynamic behavior of the system formed in PbAu collisions. These results shed some light on the dynamics of the system created in heavy-ion collisions at top SPS energy. The moderate deviations between data and the current hybrid hydrodynamics calculation could be used in future investigations to fine-tune the initial condition and transport parameters sofar rather little explained at SPS energy.
@@ Acknowledgements.
The CERES/NA45 collaboration acknowledges the good performance of the CERN PS and SPS accelerators as well as the support from the EST division. We are grateful for excellent support by the CERN IT division for the central data recording and data proccesing. This work was supported by GSI, Darmstadt, the German BMBF, the German VH-VI 146, the US DoE, the Israeli Science Foundation, and the MINERVA Foundation. We acknowledge the support by the Ministry of Education, Science and Technological Development of the Republic of Serbia throughout the theme 0102202. Z.P. acknowledges support by the State of Hesse within the Research Cluster ELEMENTS (Project ID 500/10.006). Computational resources for the SMASH-vHLLE hybrid modelling have been provided by the GreenCube at GSI. We thank Iurii Karpenko for enlightening discussions with him, and appreciate his help concerning the SMASH-vHLLE hybrid model.